# Spatio-temporal Fourier Transformer (StFT) for Long-term Dynamics Prediction

### v1

@@ Abstract
Simulating the long-term dynamics of multi-scale and multi-physics systems poses a significant challenge in understanding complex phenomena across science and engineering. The complexity arises from the intricate interactions between scales and the interplay of diverse physical processes. Neural operators have emerged as promising models for predicting such dynamics due to their flexibility and computational efficiency. However, they often fail to effectively capture multi-scale interactions or quantify the uncertainties inherent in the predictions. These limitations lead to rapid error accumulation, particularly in long-term forecasting of systems characterized by complex and coupled dynamics. To address these challenges, we propose a spatio-temporal Fourier transformer (StFT), in which each transformer block is designed to learn dynamics at a specific scale. By leveraging a structured hierarchy of StFT blocks, the model explicitly captures dynamics across both macro- and micro- spatial scales. Furthermore, a generative residual correction mechanism is integrated to estimate and mitigate predictive uncertainties, enhancing both the accuracy and reliability of long-term forecasts. Evaluations conducted on three benchmark datasets (plasma, fluid, and atmospheric dynamics) demonstrate the advantages of our approach over state-of-the-art ML methods.
@@ 1 Introduction
Predicting long-term spatio-temporal dynamics governed by partial differential equations (PDEs) is a cornerstone of scientific and engineering research, with broad applications in fields such as climate modeling, plasma science, fluid dynamics, and beyond. Traditional approaches rely heavily on numerical solvers, which discretize the domain and iteratively solve PDEs using methods like finite difference, finite element, and spectral methods (Tadmor, 2012). While effective in many scenarios, these techniques face significant limitations when applied to complex and large-scale simulations. They require substantial computational resources and exhibit poor scalability with increasing problem size, rendering them impractical for high-dimensional, large-scale, or long-term physics systems due to excessive computational costs and memory demands.
@@
Recent advances in deep learning have revolutionized the field of PDE modeling by introducing data-driven methodologies that significantly accelerate computations for science while maintaining high accuracy. Supported by universal approximation theorem (Chen & Chen, 1995), neural operators that learn the mapping between two function spaces have demonstrated great success in simulating various PDE systems across multiple scientific disciplines without retraining for new conditions (Li et al., 2020; Lu et al., 2021). With the success of transformer in natural language processing and computer vision (Vaswani, 2017; Dosovitskiy, 2020), transformer-based neural operators can process multiple input functions while enabling arbitrary querying of output function locations, offering enhanced flexibility in handling complex functional mappings (Hao et al., 2023; Li et al., 2022). A series of neural operators have been developed to address complex scientific problems, including applications in weather forecasting, turbulent fluid dynamics, and boiling phenomena (Pathak et al., 2022; Li et al., 2023b; Bi et al., 2023; Hassan et al., 2023; Lin et al., 2021).
@@
Despite success of these methods, accurately predicting the long-term evolution of these systems presents unique challenges in terms of both computational efficiency and long-term stability. The inherent multi-scale nature and multi-physics complexity of such systems necessitate methodologies that can efficiently represent and integrate dynamics across disparate spatial and temporal scales while simultaneously capturing the complex interactions between distinct physical processes, such as the influence of micro-scale turbulence on macro-scale flow (Peters, 2009; Natrajan et al., 2007). For example, large-scale atmospheric pressure systems, such as high-pressure ridges and low-pressure troughs, play a crucial role in shaping local weather patterns. Inaccurate representation of these systems can cause significant errors in forecasting rainfall, wind speed, and temperature (Wang et al., 2006; Barlow et al., 2019). In magnetically confined plasmas, the magnetohydrodynamic (MHD) instabilities caused by current or pressure gradients can limit burning plasma performance and threaten fusion device integrity (von Goeler et al., 1974; Graves et al., 2012; Seo et al., 2024). Although neural operators present advantages over traditional approaches, they still encounter challenges associated with computational complexity and the demands for scientific fidelity, especially when the underlying physics involve rapid changes or high-frequency components. These challenges are further intensified in high-resolution simulations of multi-scale scenarios. Recent efforts to address these limitations include decomposing learnable parameters (Kossaifi et al., 2023), and employing linear attention to reduce computational time  (Hao et al., 2023; Cao, 2021). Furthermore, integrating uncertainty quantification (UQ) into modeling frameworks is essential for assessing the confidence and reliability of predictions, particularly in complex systems such as in fluids and weather forecasting  (Cheung et al., 2011; Scher & Messori, 2018). However, most existing neural operators do not incorporate UQ, which is particularly important for modeling long-term dynamics, where even small errors can propagate across scales and result in significant inaccuracies.
@@
On the other hand, existing approaches for predicting spatio-temporal dynamics can be classified into two primary categories. The first category comprises models that directly forecast future states at a predetermined number of time steps based on historical observed states (Wang et al., 2024; Kontolati et al., 2024). The second category includes models that utilize an autoregressive manner, which addresses challenges of scaling and fitting complexities as a continuous-time emulator (Pathak et al., 2022; Rühling Cachay et al., 2024; Lippe et al., 2023; McCabe et al., 2023). Nevertheless, prediction errors that arise in short-term accumulate over time, resulting in unstable and inaccurate long-term forecasts. Previous work has sought to alleviate these issues through techniques such as the pushforward trick, invariance preservation, and iterative refinement (McCabe et al., 2023; Lippe et al., 2023; Brandstetter et al., 2022). However, developing robust, multi-scale modeling frameworks for long-term dynamic predictions of these systems remains crucial for improving prediction accuracy and understanding the complex interplay between scales.
@@
In this work, we introduce a spatio-temporal Fourier transformer StFT for the forecasting of multi-scale and multi-physics systems. Each StFT block captures the underlying dynamics of the physical processes at one specific scale. Through a hierarchical composition of StFT blocks across multiple scales, augmented by a generative residual correction block, our model is capable of learning the intricate interactions both within the same scales and across different scales, while also providing meaningful uncertainty quantification. The cascading StFT blocks enable our model to predict high-resolution dynamics across a spectrum of varying scales for correlated physical states. By integrating StFT within an auto-regressive framework, our method achieves superior accuracy in both short-term and long-term predictions compared to existing state-of-the-art autoregressive baselines.
Our contributions can be summarized as follows:
  • We propose Spatio-temporal Fourier transformer (StFT), a novel ML model that learns underlying dynamics across various scales while incorporating interactions among diverse physical processes.
@@
  • We propose StFT-F, which incorporates a residual correction mechanism to refine the forecasting of StFT and capture intrinsic uncertainties.
  • In each StFT block, we propose an overlapping tokenizer and a detokenizer that share boundaries between adjacent regions, enhancing the smoothness of vision representation.
  • We demonstrate the effectiveness of StFT in an autoregressive framework on a diverse set of applications. Evaluating performance across variables in each complex system, StFT outperforms the best baseline, achieving over $600\%,20\%,$ and $50\%$ higher accuracy in the plasma, fluid, and atmospheric dynamics datasets, respectively. Compared with StFT, StFT-F has an average of $5$ percent improvement while simultaneously providing meaningful uncertainty quantification.
@@ 2 Related Work
Neural Operators. There have been many popular neural operators and their variants including Fourier neural operators (Li et al., 2020; Gupta et al., 2021; Tran et al., 2021; Cao et al., 2023; Li et al., 2023a; Rahman et al., 2022), DeepONet (Lu et al., 2021; Wang et al., 2021; Jin et al., 2022; Wang et al., 2022; Kontolati et al., 2023; Prasthofer et al., 2022), transformer based operators (Hao et al., 2023; Cao, 2021; Li et al., 2022), image-to-image inspired operators (Gupta & Brandstetter, 2022; Rühling Cachay et al., 2024; Long et al., 2024). U-Net, a fundamentally hierarchical structure model, has inspired several neural operators  (Rahman et al., 2022; Liu et al., 2022b; Gupta & Brandstetter, 2022), which allows to solve multi-scale PDEs by hierarchically aggregating of feature representations progressing from fine to coarse levels. Recent work in computer vision (Liu et al., 2021; Fan et al., 2021; Zhang et al., 2022) have introduced methods for extracting multi-scale features through hierarchical architectures. However, these hierarchical models do not explicit forecast the multi-scale structures of physical processes. In contrast, our method begins with a coarse approximation that captures large-scale, low-frequency phenomena, and incrementally refines the representation over layers to resolve finer details. These structured decompositions allow for error diagnosis, enhancing interpretability of model performance across different scales, and enable targeted improvements. In contrast, U-Net like hierarchical models lack explicit refinement mechanisms, which limits the transparency and interpretability of their multi-scale representations.
Generative Models
@@
Recent years, generative models, especially diffusion models have demonstrated huge success in various domains, including vision, audio, robotics, and medical field (Ho et al., 2020; Song et al., 2020; Tian et al., 2024; Kong et al., 2020; Wolleb et al., 2022). Video generation techniques are closely related to spatio-temporal dynamics prediction, as demonstrated in (Ho et al., 2022; Voleti et al., 2022; Singer et al., 2022). Recent work emulates the forward and inverse transformations of diffusion models to learn the spatio-temporal dynamics (Rühling Cachay et al., 2024). As an alternative approach in generative modeling, flow matching has been introduced to enable fast sampling and has since been applied to video generation (Lipman et al., 2022; Liu et al., 2022a; Albergo & Vanden-Eijnden, 2022; Polyak et al., 2024; Esser et al., ). While video generation typically explores a range of creative and diverse possibilities from text or image prompts, forecasting spatial-temporal dynamics driven by PDEs necessitates more than mere statistical resemblance - it requires each prediction is firmly grounded in the underlying physical laws. To achieve accurate forecasting while capturing the inherent stochasticity of physical processes, our work incorporates a flow matching block following the proceeding StFT blocks. This approach allows our model to precisely recover the true dynamics while simultaneously offering meaningful uncertainty quantification.
@@ 3 Method
We consider an autoregressive formulation for long-term multi-scale spatiotemporal physical processes. We define a vector $\tilde{u}_{t}$ representing the historical snapshots of the multi-physics variables at timestamps from $t-k+1$ to $t$, in a total of $k$ snapshots. Specifically, $\tilde{u}_{t}=[u_{t},u_{t-1},\dots,u_{t-k+1}]$. We formulate the probabilistic one-step forward neural operator StFT-F as
$u_{t}$ $=\mathcal{F}_{\theta_{d}}(\tilde{u}_{t-1})+r_{t},$ (1)
$r_{t}$ $\sim\mathcal{P}_{\theta_{g}}(r|\tilde{u}_{t-1},\mathcal{F}_{\theta_{d}}(\tilde{u}_{t-1})),$ (2)
where $\mathcal{F}_{\theta_{d}}$ denote the StFT operator, a deterministic forecasting parameterized by $\theta_{d}$, and $\mathcal{P}_{\theta_{g}}$ is the generative flow matching block parameterized by $\theta_{g}$ for refining the forecasting of StFT and providing uncertainty quantification. $\mathcal{F}_{\theta_{d}}$ represents the deterministic evolution of the system that encapsulates the predictive dynamics. The term $\mathcal{P}_{\theta_{g}}(r|\tilde{u}_{t-1},\mathcal{F}_{\theta_{d}}(\tilde{u}_{t-1}))$ captures the probabilistic nature of the residual of the model. It represents a distribution conditioned on the current state $\tilde{u}_{t-1}$, and the deterministic prediction, modeling the uncertainty or variations that are not captured by the deterministic component. The residual $r_{t}$ calculates deviations from the deterministic prediction, and its distribution allows the model to account for noise or inherent stochasticity in the physical processes. Therefore, by sampling residual $r_{t}$, our model learns stochastic trajectories from data. Besides providing predictive uncertainties caused by inherent noise, these stochastic trajectories can help study the long-term behavior, stabilities, and bifurcations in stochastic systems.
Figure 1 presents the overview of the StFT’s architecture. In the following subsections, we introduce the overlapping tokenizer, detokenizer, StFT block, and the residual refinement mechanism based on flow matching.

### v2

# StFT: Spatio-temporal Fourier Transformer for Long-term Dynamics Prediction

@@ Abstract
Simulating the long-term dynamics of multi-scale and multi-physics systems poses a significant challenge in understanding complex phenomena across science and engineering. The complexity arises from the intricate interactions between scales and the interplay of diverse physical processes, which manifest in PDEs through coupled, nonlinear terms that govern the evolution of multiple physical fields across scales. Neural operators have shown potential in short-term prediction of such complex spatio-temporal dynamics; however, achieving stable high-fidelity predictions and providing robust uncertainty quantification over extended time horizons remains an open and unsolved area of research. These limitations often lead to stability degradation with rapid error accumulation, particularly in long-term forecasting of systems characterized by multi-scale behaviors involving dynamics of different orders. To address these challenges, we propose an autoregressive Spatio-temporal Fourier Transformer (StFT), in which each transformer block is designed to learn the system dynamics at a distinct scale through a dual-path architecture that integrates frequency-domain and spatio-temporal representations. By leveraging a structured hierarchy of StFT blocks, the resulting model explicitly captures the underlying dynamics across both macro- and micro- spatial scales. Furthermore, a generative residual correction mechanism is introduced to learn a probabilistic refinement temporally while simultaneously quantifying prediction uncertainties, enhancing both the accuracy and reliability of long-term probabilistic forecasting. Evaluations conducted on three benchmark datasets (plasma, fluid, and atmospheric dynamics) demonstrate the advantages of our approach over state-of-the-art ML methods.
footnotetext: Corresponding author (zhebai@lbl.gov)
[…]
Spatio-temporal modeling, long-range forecasting, multi-scale physical systems, hierarchical scale learning, flow matching, uncertainty quantification.
@@ 1 Introduction
Predicting the long-term spatio-temporal dynamics of systems governed by partial differential equations (PDEs) is a cornerstone of scientific and engineering research, with broad applications in fields such as earth system modeling, plasma science, fluid dynamics, and beyond. Traditional approaches rely heavily on numerical solvers, which discretize the domain and iteratively solve PDEs using methods including finite element, finite volume and spectral methods (tadmor2012review, ). While effective in many scenarios, these techniques face limitations when applied to multiphysics systems characterized by complex dynamics and multiscale behaviors. They require substantial computational resources and exhibit poor scalability with increasing problem size, rendering them impractical for high-dimensional, large-scale, or long-term physics systems due to excessive computational costs and memory demands.
@@
Recent advances in deep learning have revolutionized the field of PDE modeling by introducing data-driven methodologies that significantly accelerate computations for science while maintaining high accuracy. Inspired by the universal approximation theorem (chen1995universal, ), neural operators that learn the mapping between two function spaces have demonstrated great success in simulating various PDE systems across multiple scientific disciplines without retraining for new conditions (li2020fourier, ; lu2021learning, ). Building on the success of transformers in natural language processing and computer vision (vaswani2017attention, ; dosovitskiy2020image, ), transformer-based neural operators process multiple input functions while enabling arbitrary querying of output function locations, offering enhanced flexibility in handling complex functional mappings (hao2023gnot, ; li2022transformer, ). A series of neural operators have been developed to address complex scientific problems, including weather forecasting, turbulent fluid dynamics, and boiling phenomena (pathak2022fourcastnet, ; li2023long, ; bi2023accurate, ; hassan2023bubbleml, ; lin2021operator, ).
@@
Despite the success of these methods, accurate and long-term predictions of complex physical systems remain challenging, primarily due to the requirements for numerical stability, high-fidelity modeling, and reliable uncertainty quantification over extended horizons. The inherent multi-scale nature and multi-physics complexity of such systems necessitate methodologies that can efficiently represent and integrate dynamics across disparate spatial and temporal scales while simultaneously capturing the complex interactions between distinct physical processes, such as the influence of micro-scale turbulence on macro-scale flow in fluids and combustion (peters2009multiscale, ; natrajan2007statistical, ). For large-scale atmospheric pressure systems, high-pressure ridges and low-pressure troughs play a crucial role in shaping local weather patterns; inaccurate representation of those structures can cause significant errors in forecasting rainfall, wind speed, and temperature (wang2006large, ; barlow2019north, ). In magnetically confined plasmas, multiphysics arises from the coupling of physical processes that govern plasma behavior, including electromagnetic fields, turbulence transport, thermodynamics, and particle interactions that are potentially coupled with kinetic models in high-fidelity simulations. The magnetohydrodynamic (MHD) instabilities caused by current or pressure gradients can limit burning plasma performance, and threaten fusion device integrity (von1974studies, ; graves2012control, ; seo2024avoiding, ). Furthermore, integrating uncertainty quantification (UQ) into modeling frameworks is essential for assessing the confidence and reliability of predictions in such complex systems (cheung2011bayesian, ; scher2018predicting, ; kruger2024thinking, ). Although neural operators present advantages over traditional approaches, they still encounter challenges associated with the demands for scientific fidelity and stability, especially when the underlying physics exhibit rapid changes or high-frequency components. These issues are further intensified in high-resolution simulations of multi-scale scenarios. Recent efforts to address these limitations include $P^{2}C^{2}$Net, which encodes a high-order numerical scheme with boundary condition encoding into neural networks (NEURIPS2024_7f605d59, ), and Dyffusion, which trains a forecasting network and an interpolation network that allows for continuous time sampling and multi-step prediction for long-range forecasting (ruhling2023dyffusion, ). However, most existing neural operators lack built-in mechanisms for uncertainty quantification, which is particularly critical for reliable modeling of long-term dynamics, where even small errors can propagate across scales and result in significant inaccuracies.
@@
Existing approaches for predicting spatio-temporal dynamics can be broadly be classified into two primary categories. The first category comprises models that directly forecast future states at fixed time horizons using a sequence of past observations (wang2024bridging, ; kontolati2024learning, ). The second category includes models that utilize an autoregressive manner, which addresses challenges of scaling and fitting complexities as a continuous-time emulator (pathak2022fourcastnet, ; ruhling2023dyffusion, ; lippe2023pde, ; mccabe2023towards, ). Generally, prediction errors incurred in the short term can accumulate, leading to instability and reduced accuracy in long-term forecasts. To mitigate these issues, previous work has proposed techniques such as the pushforward trick, invariance preservation, and iterative refinement (mccabe2023towards, ; lippe2023pde, ; brandstetter2022message, ). Nevertheless, the development of multi-scale modeling frameworks for long-term dynamic prediction remains crucial for capturing the interactions across scales and enhancing prediction accuracy. Concurrently, incorporating uncertainty quantification is critical for identifying the spatial and temporal regions where predictive confidence deteriorates, especially in complex systems where localized uncertainties can influence global dynamics over time.
@@
In this work, we introduce an autoregressive Spatio-temporal Fourier Transformer (StFT), for long-range forecasting of multi-scale and multi-physics systems. At each level of the spatial hierarchy, one StFT block models the physical dynamics associated with a distinct spatial scale or receptive field, as inferred from the spatiotemporal data. Each StFT block adopts a dual-path architecture: (1) the frequency path captures large-scale dynamics by operating in the Fourier domain, focusing on low-frequency components that are critical for modeling global behavior. (2) the spatio-temporal path operates in the full physical space incorporating all spatio-temporal features to capture fine-scale features. Through a hierarchical composition of StFT blocks across multiple scales, augmented by a generative residual correction block, the resulting model learns the intricate interactions both within the same scales and across different scales. Moreover, it produces uncertainty estimates at each spatial and temporal point, enabling assessment of prediction confidence throughout the forecast. The cascading StFT blocks enable our model to predict high-resolution dynamics across a spectrum of varying scales in correlated physical processes. By integrating StFT within an auto-regressive framework, our method achieves superior accuracy in long-term predictions compared to existing state-of-the-art autoregressive baselines. Our contributions are summarized as follows:
  • We propose Spatio-temporal Fourier transformer (StFT), a novel ML model that learns underlying dynamics across spatial scales for multi-physics systems via a dual-path (frequency and spatio-temporal path) architecture, which effectively captures both the global, large-scale structures and local, fine-scale features.
@@
  • We propose StFT-F, which incorporates a probabilistic residual correction mechanism to refine the forecasting of StFT temporally and provide pointwise uncertainty quantification.
  • We propose an overlapping tokenizer and a detokenizer that share regions between adjacent patches, improving spatial smoothing and reducing discontinuity artifacts.
  • We demonstrate the effectiveness of StFT in an autoregressive framework on a diverse set of applications including the plasma, fluid, and atmospheric dynamics. Evaluating performance across variables, StFT outperforms the best baselines. Its probabilistic variant, StFT-F further improves average forecasting accuracy by 5% and produces uncertainty estimates that are empirically calibrated, as demonstrated through confidence-based evaluation.
@@ 2 Related Work
Neural Operators. Neural operator architectures and their variants have been proposed, including Fourier neural operators (li2020fourier, ; gupta2021multiwavelet, ; tran2021factorized, ; cao2023lno, ; li2023fourier, ; rahman2022u, ), DeepONet (lu2021learning, ; wang2021learning, ; jin2022mionet, ; wang2022improved, ; kontolati2024learning, ; prasthofer2022variable, ), transformer based operators (hao2023gnot, ; cao2021choose, ; li2022transformer, ), and image-to-image operators (gupta2022towards, ; long2024arbitrarily, ). U-Net, a fundamentally hierarchical structure model, which has inspired several neural operators  (rahman2022u, ; liu2022ht, ; gupta2022towards, ), allows solutions to multi-scale PDEs by hierarchically aggregating feature representations of progressively coarser spatial resolutions. Recent work in computer vision (liu2021swin, ; fan2021multiscale, ; zhang2022nested, ) have introduced methods for extracting multi-scale features through hierarchical architectures. However, these hierarchical models do not explicitly model or forecast the multi-scale structures of physical processes, which limit the transparency and interpretability of their representations across scales. In contrast, our method begins with a coarse approximation that captures large-scale, low-frequency phenomena, and incrementally refines the representation over layers to resolve finer details. These structured decompositions enable error diagnosis, enhance interpretability of model performance of different scales, and allow for targeted improvements with an explicit refinement mechanism.
@@
Generative Models. Generative models, especially diffusion models have demonstrated success in various domains, including vision, audio, robotics (ho2020denoising, ; song2020score, ; tian2024transfusion, ; kong2020diffwave, ; wolleb2022diffusion, ), and relevance to spatio-temporal dynamics prediction (ho2022video, ; voleti2022mcvd, ; singer2022make, ; ruhling2023dyffusion, ). As an alternative approach in generative modeling, flow matching has been introduced to support efficient sampling and has since been applied to video generation (lipman2022flow, ; liu2022flow, ; albergo2022building, ; polyak2024movie, ; esser2403scaling, ). While video generation typically explores a range of creative and diverse possibilities from text or image prompts, forecasting spatial-temporal dynamics driven by PDEs necessitates more than mere statistical resemblance - it requires each prediction is firmly grounded in the underlying physics. To achieve accurate forecasting while capturing the inherent stochasticity of physical processes, our work incorporates a flow matching block following the proceeding StFT blocks. This enables the model to align its prediction distribution with the underlying physical dynamics and generate calibrated uncertainty estimates via confidence-based metrics.
@@ 3 Method
Figure 1: Left: Overview of the proposed StFT and StFT-F model. The model predicts $u_{t+1}$ using the past $k$ snapshots $\tilde{u}_{t}$, employing $L$ spatial refinements from coarse to fine scales through the proposed Spatiotemporal Fourier Transformer (StFT) blocks. Bottom left: an illustration of the overlapping tokenizer, where the patch size is $2\times 2$, and the overlapping number is $1\times 1$. Right: Illustration of the proposed StFT block. First, the past snapshots $\tilde{u}_{t}$ and a coarser prediction from the previous layer are temporally stacked. The stacked discretizations are passed to the overlapping tokenizer to generate tokens for each variable. Next, tokens corresponding to different variables in the same spatial are mixed through a variable mixer. Two paths of transformation, frequency path and spatiotemporal path, process the frequency embeddings and the spatiotemporal embeddings respectively. The finer prediction for timestamp $t+1$ is obtained after passing through the overlapping detokenizer.
Formulation. We consider an autoregressive formulation for long-term multi-scale spatiotemporal physical processes. We define a vector $\tilde{u}_{t}$ representing the historical snapshots of the multi-physics variables at timestamps from tk+1t-k+1 to $t$, in a total of $k$ snapshots of $\tilde{u}_{t}=[u_{t},u_{t-1},\dots,u_{t-k+1}]$ specifically. We formulate the probabilistic one-step forward neural operator StFT-F as
$u_{t+1}=\mathcal{F}_{\theta_{d}}(\tilde{u}_{t})+r_{t+1},r_{t+1}\sim\mathcal{P}_{\theta_{g}}(r|\tilde{u}_{t},\mathcal{F}_{\theta_{d}}(\tilde{u}_{t})),$ (1)
where $\mathcal{F}_{\theta_{d}}$ denote the StFT operator, a deterministic forecasting parameterized by $\theta_{d}$, and $\mathcal{P}_{\theta_{g}}$ is the generative flow matching block parameterized by $\theta_{g}$ for refining the forecasting of StFT while quantifying uncertainty estimates. $\mathcal{F}_{\theta_{d}}$ represents the deterministic evolution of the system that encapsulates the dynamics from multi-scale spatial refinement in StFT. The residual refinement $\mathcal{P}_{\theta_{g}}(r|\tilde{u}_{t},\mathcal{F}_{\theta_{d}}(\tilde{u}_{t}))$ captures the probabilistic nature of the residual from the generative model. It represents a prediction distribution conditioned on the current state $\tilde{u}_{t}$ and the deterministic prediction $\mathcal{F}_{\theta_{d}}(\tilde{u}_{t})$, modeling the uncertainty or variations that missed by the deterministic component. The residual $r_{t+1}$ calculates deviations from the deterministic prediction, and its distribution allows the model to account for noise or inherent stochasticity in the physical processes. Therefore, by sampling residual $r_{t}$, our model learns stochastic trajectories from data. Besides providing prediction uncertainties, these stochastic trajectories can help study the long-term behavior, stabilities, and bifurcations in stochastic systems  (lucor2003predictability, ; kramer2022probabilistic, ).
Figure 1 presents the overview of the StFT-F, the overlapping tokenizer, and the design of the StFT block. Algorithm  1 and  2 detail the design of our model. The following subsections introduce StFT with the overlapping tokenizer/detokenizer and the residual correction mechanism based on flow matching.